OpenStudy (oraclethinktank):

How many solutions exist for the system below? In two or more complete sentences, give the solution and explain how you solved the system. y = -2x + 3 y = -2x * 3^x

OpenStudy (oraclethinktank):

I already know that there is no solution. I just do not know the equation to solve the system.

OpenStudy (photon336):

this right? \[y = -2x+3 \] \[y = -2x*3^{x}\]

OpenStudy (oraclethinktank):

The first one is correct, the second one, there is no x next to the 2. That's a mistake on my part. It would be y = -2*3^x

OpenStudy (photon336):

\[y = -2*3^{x}\]

OpenStudy (oraclethinktank):

Yeah.

OpenStudy (photon336):

hmm solve the top equation for x for me y = -2x+3

OpenStudy (oraclethinktank):

By graphing, the slope is -2 and the y intercept is 3

OpenStudy (photon336):

Yeah taken alone that is correct.

OpenStudy (photon336):

y = -2x+3 y = -2*3^x \[-2x+3 = -2*3^{x}\] probably the best thing to do if you want to find the intersection of two equations is to set them equal to each-other.

OpenStudy (oraclethinktank):

The question already has a graph shown. There are no intersection points, Therefore there is no solution.

OpenStudy (photon336):

there is no graph It's not posted.

OpenStudy (photon336):

if they gave you a graph you can easily say that no intersection points exist therefore there is no solution to your equation.

OpenStudy (oraclethinktank):

Not entirely, I'm supposed to convert it to the Quadratic formula to show that there is no solution. I wish it was as easy as stating that there is no solution.

OpenStudy (photon336):

\[y = -2*3^{x}\] that's exponential though.. don't know how they could ask for that. you sure this is the right question?

OpenStudy (oraclethinktank):

Yes. That is exactly how it is shown. I don't understand it myself.

OpenStudy (oraclethinktank):

This is the graph shown.

OpenStudy (oraclethinktank):

The system of equations makes this graph.

OpenStudy (photon336):

yep no solution to that.

OpenStudy (photon336):

well, you can't solve anything because there isn't an intersection point so there's no solution to it but if you were to solve it you could set the two equal to each-other.

OpenStudy (oraclethinktank):

So by setting the two equal to each-other, the next step would be to substitute them into the quadratic formula.

OpenStudy (oraclethinktank):

How would I do that?

OpenStudy (oraclethinktank):

@Photon336

OpenStudy (photon336):

@oraclethinktank why the quadratic formula?